# Multi-task learning for molecular electronic structure approaching coupled-cluster accuracy

### v1

@@ II Results
To test the generalization capability and data efficiency of our model, we train the model with various training dataset size $N_{\rm train}$, ranging from 10 to 7440 atomic configurations. The testing standard deviation error of different trained properties exhibit a decreasing trend when the training dataset size increases, as shown in Fig. S1a, indicating effective model generalization. Notably, the energy error has the fastest drops with a slope of -0.38 (that means the testing error $\propto N_{\rm train}^{\rm-0.38}$). In comparison, some of the recently developed advanced machine learning potential (that directly learn energies and their derivatives) exhibits a lower slope of about -0.25 [7, 9]. This implies potential advantage of our multi-task method: as our multi-task method learns different molecular properties through a shared representation (the electronic structure), the domain information learned from one property can help the model’s generalization on predicting other properties [39], providing outstanding data efficiency.
@@
Then, we benchmark the computational costs and prediction accuracy of our model trained on 7440 atomic configurations with 70 different molecules, which will be used in the rest of this paper. Our EGNN method exhibits significantly smaller computational cost and slower scaling with system size, as compared with both the B3LYP method and CCSD(T) method (Fig. S1b). Compared to the B3LYP functional, our method avoids the expensive calculation of the exact exchange thus requiring much smaller computational costs [27]. Using the gold-standard CCSD(T) calculation as a reference, the prediction accuracy of our EGNN method on various molecular properties is compared with that of the B3LYP method, as shown in Fig. S1c and Table 1. The EGNN predictions on ground-state properties $E$, $\vec{p}$, $\mathbf{Q}$, $C_{i}$, and $B_{ij}$ exhibits significantly smaller error than the B3LYP method for all molecules in both the in-domain (ID) and out-of-domain (OOD) testing dataset. Remarkably, the root-mean-square error of the ground-state energies predicted by the EGNN is about 0.1 kcal/mol (about 4 meV) per atom in both the ID and OOD dataset. Therefore, the model predictions on reaction energies can approach the quantum chemical accuracy (assuming that on average 1 mol molecules in reactants contain $\sim$10 mol atoms). The EGNN predictions on excited-state properties $E_{\rm g}$ and $\alpha$ also exhibit better overall accuracy than the B3LYP method in both the ID and OOD testing dataset (Table 1), though there are certain molecules where the B3LYP performs better. The statistical distribution of the EGNN and B3LYP’s prediction error are shown in Fig. S1 in SI. Although the B3LYP gives smaller median error for $\alpha$, the EGNN is less likely to make large errors (more robust), leading to better root-mean-square error.
@@
Hydrocarbon molecules have a vast structural space, including various types of local atomic environments (see a few examples below Fig. S1c). Our EGNN method provides a single model that exhibits generalization capability among the vast hydrocarbon structural space. To examine the model’s generalization to more complex structures, we apply the EGNN model to a series of aromatic hydrocarbon molecules synthesized in experiments [35]. The gas phase standard enthalpy of formation $H_{\rm f}$ is an essential thermochemical property of molecules that can be accurately measured. In this regard, we use the EGNN model to predict $H_{\rm f}$ of various aromatic molecules in a comprehensive experimental review paper Ref. [35]. Among the 71 molecules lists in Ref. [35], we calculate $H_{\rm f}$ of those molecules with serial numbers (defined in Ref. [35]) dividable by 4 and compare them with experiments, as shown in Fig. 3. The selected molecules cover various classes of aromatic molecules, including polycyclic aromatic hydrocarbons, cyclophanes, polyphenyl, and nonalternant hydrocarbons. The EGNN predictions on $H_{\rm f}$ are well consistent with experiments on all molecules, and their difference is only around 0.1$\sim$0.2 kcal/mol per atom. Note that the EGNN prediction error is on the same order of magnitude as the experimental error bar (though numerically larger), indicating high prediction accuracy.
@@
Figure 4: EGNN predictions for semiconducting polymers. (a) Atomic structure and HOMO wavefunctions of t-PA, polyphenylene PPP, and c-PA. The HOMO wavefunctions are visualized by isosurface at the level of ±$\pm 0.01$-2/3 (positive isosurface colored blue and negative isosurface colored yellow). (b) Energy gap and (c) static electric polarizability of t-PA (blue lines), PPP (green lines), and c-PA (orange lines) with different polymer chain length. Longitudinal polarizability $\alpha_{xx}$, horizontal polarizability $\alpha_{yy}$, and vertical polarizability $\alpha_{zz}$ are shown as silid, dashed, and dotted lines, respectively. Squares (blue for t-PA and green for PPP) represent literature values for polymers in experiments [40, 41] and correlated calculations [42], and blue dots represent literature values for t-PA oligomers from the MP2 correlated calculations [43].
@@
Finally, we apply the EGNN model to semiconducting polymers consisting of hundreds of atoms, which are difficult to calculate by rigorous correlated methods such as CCSD(T). Semiconducting polymers are organic macromolecules with small bandgap and high electrical conductivity compared to insulating polymers. Because of these electronic features, semiconducting polymers attract broad research interests both for the fundamental understanding of quantum chemistry [41] and for applications in semiconductor industry [44]. The essential electronic properties of semiconducting polymers originates from the $\pi$-bonds with delocalized molecular orbitals. As the delocalized molecular orbitals extend through the whole macroscopic molecule (Fig. 4a), the polymers’ electronic properties also involve long-range correlation, making it challenging for ML methods. Therefore, it is important to examine whether the EGNN model can capture semiconducting polymers’ electronic properties involving delocalized molecular orbitals.
@@
Three kinds of semiconducting polymers, trans-polyacetylene (t-PA), cyclic polyacetylene (c-PA), and polyphenylene (PPP), are studied using our EGNN model. The model correctly captures the delocalized $\pi$-bond feature of frontier orbitals (highest occupied molecular orbital (HOMO) and lowest unoccupied molecular orbital (LUMO)), as shown in Fig. 4a. In the t-PA, the HOMO and LUMO are on the carbon-carbon double and single bonds, respectively, consistent with the predictions from the renowed Su–Schrieffer–Heeger (SSH) model [41]. In the c-PA, the left/right mirror symmetry of the molecule is correctly reflected in the predicted frontier orbitals, which exhibits odd parity upon mirror reflection. Both the HOMO and LUMO have two anti-phase boundaries where the wavefunctions shift from double bonds to single bonds. The same symmetry is also reflected in the PPP with odd-parity HOMO and even-parity LUMO. It should be emphasized that the molecular orbitals are not included as fitting targets in our model training. However, as the model learned relevant properties such as electric moments, atomic charge, and bond order, it gives at least qualitatively correct predictions for frontier orbitals in these semiconducting polymers.
@@
Various important electronic properties of semiconducting polymers depend on the chain length, including bandgap $E_{\rm g}$ and polarizability $\alpha$. We calculate such chain-length dependence (up to more than 400 atoms) using our EGNN method, as shown in Fig. 4b,c. One can see that $E_{\rm g}$ is larger for oligomers with shorter chain length due to the quantum confinement effect and converges to a smaller value for long polymer. The converged bandgap for long t-PA and PPP polymers calculated by our EGNN are in reasonable agreement with the experimental values [41, 40], as shown as squares in Fig. 4b. The longitudinal static electric polarizability $\alpha_{xx}$ (per monomer) is positively related to the polymer chain length. This is because in longer chains, more delocalized electron distributions can have larger displacements under external electric field. The predicted $\alpha_{xx}$ for t-PA oligomers and PPP polymers are in perfect agreement with previous correlated calculations using the high-accuracy MP2 method [42, 43], as shown in Fig. 4c. The chain length-dependent $E_{\rm g}$ and $\alpha$ of cyclic PA, to the best of our knowledge, has not been reported. We provide their values as a prediction to be examined by future work.
@@ III Conclusion and Outlook
In this work, we developed a multi-task learning method for predicting molecular electronic structure with coupled-cluster-level accuracy. In our method, an EGNN is trained on a series of molecular properties in order to capture their shared underlying representation, the effective electronic Hamiltonian. Our EGNN model shows lower computational costs and higher prediction accuracy on various molecular properties than the widely-used B3LYP functional.
Compared to other works that take the electronic structure as a direct fitting target [15, 23, 21, 22], our method takes the electronic structure as a shared representation to predict various molecular properties. This approach, on the one hand, enables training on data beyond the DFT accuracy; on the other hand, relieve the burden of fitting all matrix elements of the electronic Hamiltonian, which are not direct quantum observable. The later enables us to go beyond the minimal basis used in previous works by a relatively small neural-network with only 511,589 parameters. With the electronic structure as a physics-informed representation, our method exhibits generalization capability for challenging systems with delocalized molecular orbitals. In such systems, atomic structure at one position has long-range influence to electronic features (such as electron density) at another distant position. Although the direct output of our EGNN, the matrix elements of $V_{\theta}$, only depends on local atomic environments, the long-range influence is captured through the output-layer calculations based on the rules of quantum mechanics. In comparison, such long-range influence is unlikely to be captured by directly using GNN or local atomic descriptors to predict molecular properties without a physics-informed representation.
Although the current EGNN model is only applicable to hydrocarbon molecules, our method is generally applicable to systems with different elements. Applying the method to datasets with more elements can produce a general-purpose electronic structure predictor, which is left to future work.
@@ IV.1 Graph encoding of atomic configuration
As the hydrocarbon molecules we study are all close-shell molecules, we use spin-restricted DFT calculations to obtain ${\bf F}$ and we assume ${\bf V}_{\theta}$ is spin-independent throughout this paper. Namely, the spin-up and spin-down molecular orbitals and energy levels are the same, and all molecular orbitals are either doubly occupied or vacant. The electronic energy of the BP86 DFT calculation $E_{\rm BP86}$ equals the band structure energy $2\sum_{i=1}^{n_{e}/2}\epsilon_{i}$ (where $\epsilon_{i}$ is the $i$th molecular orbital energy level and $n_{e}$ is the number of electrons) plus a many-body energy $E_{\rm MB}$:
@@
where ${\rm eig}_{i}$ is a function that returns the $i$th lowest eigenvalue of a matrix, and we use the fact that the energy level $\epsilon_{i}$ is the eigenvalue of the Lowdin-symmetrized Hamiltonian ${\rm eig}_{i}(S^{-1/2}FS^{-1/2})$ [31]. After this transformation, the KS effective Hamiltonian (both before and after the $V_{\theta}$ correction) already includes the many-body energy, and the total electronic energy is just the summation of band structure energies.
@@ IV.2 Architecture of the convolutional layer
In the following technical description, we use terminologies defined in the e3nn documentation. Unfamiliar readers can refer to Ref. [26] for further information. In the convolutional layer, the input feature first goes through a $N_{\rm species}\times N_{\rm species}$ linear transformation (the first Self-Interaction block) and an activation layer (the first non-Linearity block). All activation layers in our EGNN are realized by tanh function acting on scalar features.
@@
Finally, the output features are used to construct the correction matrix $V_{\theta}$ through Eq. (2). $V_{\rm node}(\mathbf{x}_{i,\rm out})$ first apply a linear layer from input dimension of Irreps2 to output dimension $\text{Irreps}(i)^{\otimes 2}$, where:
@@
The output dimension corresponds to the irreducible representation of the block diagonal terms of the Hamiltonian (as cc-pVDZ basis of hydrogen includes two s orbitals (0e) and one group of p orbitals (1o); and that of carbon includes three s orbitals (0e), two groups of p orbitals (1o), and one group of $d$ orbitals (2e)). The output is then arranged into the $N_{i}\times N_{i}$ matrix form, $V_{i,\rm out}$, according to the Wigner-Eckart theorem [22], and symmetrized to obtain $V_{\rm node}(\mathbf{x}_{i,\rm out})=\frac{\lambda_{V}}{2}(V_{i,\rm out}+V_{i,\rm out}^{T})$. $\lambda$ is a constant hyperparameter set as 0.2 for our model. Similarly, the off-diagonal term $V_{\rm edge}(\mathbf{f}_{ij,\rm out})$ in Eq. (2) applies a linear layer from input dimension of Irreps2 to output dimension Irreps($i,j$) that equals:

### v2

@@ II.2 Model Performance and Applications
Benchmark of Model Performance. The model’s generalization capability from small molecules to large molecules is essential for its usefulness on complex systems where coupled cluster calculations cannot be implemented on current computational platforms, due to their formidable computational costs. To test the generalization capability and data efficiency of our model, we train the model with varied training dataset size $N_{\rm train}$, ranging from 10 to 7440 atomic configurations. The testing root-mean-square error (RMSE, absolute error in atomic units) of different trained properties exhibit a decreasing trend when the training dataset size increases, as shown in Fig. S1a, indicating effective model generalization. Notably, the energy error has the fastest drops with a slope of $-0.38$ (that means the testing error $\propto N_{\rm train}^{\rm-0.38}$). In comparison, some of the recently developed advanced ML potentials (that directly learn energies and their derivatives) exhibits lower slopes of about -0.25 [7, 9]. This implies potential advantage of our multi-task method: as our multi-task method learns different molecular properties through a shared representation (the electronic structure), the domain information learned from one property can help the model’s generalization on predicting other properties [44], providing outstanding data efficiency.
Figure 3: Validation of the EGNN’s predictions on gas phase aromatic hydrocarbon molecules, as compared with experimental results. (a) Standard enthalpy of formation. The EGNN predictions and experimental values from Ref. [45] (right axis) are compared for 11 molecules (see SI Table II for details). The difference between the EGNN method and experimental values are shown by the orange line, and the experimental uncertainty is shown by the red line (left axis). (b) Infrared spectrum of benzene. The experimental data is from the NIST Chemistry WebBook [46]. Vibration modes corresponding to the peaks are labeled following the convention in Ref. [47].
@@
Then, we benchmark the computational costs and prediction accuracy of our model trained on 7440 atomic configurations with 70 different molecules, which will be used in the rest of this paper. Our EGNN method exhibits significantly smaller computational cost and slower scaling with system size, as compared with the hybrid functional, double hybrid functional [39], and CCSD(T) method (Fig. S1b). Compared to the hybrid functional, our method avoids the expensive calculation of the exact exchange thus requiring much smaller computational costs [31]. Using the gold-standard CCSD(T) calculation as a reference, the prediction accuracy of our EGNN method on various molecular properties is compared with that of the several theory functionals and existing ML methods, as shown in Fig. S1c and Table 1. The comparison is implemented on both an in-domain (ID) and an out-of-domain (OOD) testing dataset of hydrocarbon molecules. The EGNN predictions exhibit smaller RMSE than the hybrid, double hybrid, and DM21 [14] functional on most compared molecular properties (except electric dipole moment on the OOD dataset, where double hybrid gives the smallest RMSE). Remarkably, the RMSE of the combination energy predicted by the EGNN is about 0.1 kcal/mol (about 4 meV) per atom in both the ID and OOD dataset. Our method exhibits similar RMSE of combination energy compared to the AIQM1 ML potential that features its chemical accuracy energy predictions. These results confirm that our EGNN’s predictions on reaction energies can approach the quantum chemical accuracy (assuming that on average 1 mol molecules in reactants contain $\sim$10 mol atoms). Note that besides the ground-state properties, our EGNN also provides excited-state property $E_{\rm g}$ and linear response property $\alpha$ with better overall accuracy than other methods in comparison (Table 1), though there are certain molecules where the hybrid functional performs better. The statistical distribution of the prediction errors of the EGNN and B3LYP hybrid functional are shown in Fig. S1 in SI. Although the hybrid functional gives smaller median error for $\alpha$, the EGNN is less likely to make large errors (more robust), leading to better root-mean-square error.
@@
Aromatic Molecules. Hydrocarbon molecules have a vast structural space, including various types of local atomic environments (see a few examples below Fig. S1c). Our EGNN method provides a single model that exhibits generalization capability among the vast hydrocarbon structural space (cf. Fig. S1). To further examine the model’s generalization capability in more complex structures, we apply the EGNN model to a series of aromatic hydrocarbon molecules synthesized in experiments [45]. The gas phase standard enthalpy of formation $H_{\rm f}$ is an essential thermochemical property of molecules that can be accurately measured in experiments. In this regard, we use the EGNN model to predict $H_{\rm f}$ of various aromatic molecules in a comprehensive experimental review paper Ref. [45]. Among the 71 molecules lists in Ref. [45], we calculate $H_{\rm f}$ of those molecules with serial numbers (defined in Ref. [45]) dividable by 4 and compare them with experiments, as shown in Fig. 3a. The selected molecules cover various classes of aromatic molecules, including polycyclic aromatic hydrocarbons, cyclophanes, polyphenyl, and nonalternant hydrocarbons. The EGNN predictions on $H_{\rm f}$ are well consistent with experiments on all molecules, and their difference is only around $0.1\sim 0.2$ kcal/mol per atom. Note that the EGNN prediction error is on the same order of magnitude as the experimental error bar (though numerically larger), indicating high prediction accuracy.
Besides thermochemical properties, our EGNN model can also predict spectral properties, as shown in Fig. 3b. Especially, infrared (IR) spectrum reflects essential information of molecular vibrational modes and their interaction with light. In previous work of ML electronic structure [20], although the predicted peak positions of the IR spectrum are well consistent with the experiment, the predicted peak intensity are inconsistent with the experiment. In comparison, our EGNN model predicts both the peak positions and intensity well consistent with the experiment, providing both the fundamental bands and combination bands known as “benzene fingers” in the IR spectrum. The good consistency of peak intensity is attributed to accurate predictions on the transition dipole moments that determine the intensity of light-matter interaction. Details on calculating the IR spectrum is elaborated in SI section S3.
@@
Figure 4: EGNN predictions for the electronic proerpties of semiconducting polymers. (a) Atomic structure and HOMO wavefunctions of t-PA, polyphenylene PPP, and c-PA. The HOMO wavefunctions are visualized by isosurfaces at the level of ±$\pm 0.01$-2/3 (positive isosurface colored blue and negative isosurface colored yellow). (b) Energy gap and (c) static electric polarizability of t-PA (blue lines), PPP (green lines), and c-PA (orange lines) with different polymer chain length. Longitudinal polarizability $\alpha_{xx}$, horizontal polarizability $\alpha_{yy}$, and vertical polarizability $\alpha_{zz}$ are shown as solid, dashed, and dotted lines, respectively. Squares (blue for t-PA and green for PPP) represent literature values for polymers in experiments [48, 49] and correlated calculations [50], and blue dots represent literature values for t-PA oligomers from the MP2 correlated calculations [51].
@@
Large-scale Semiconducting Polymers. Besides small molecules, we also apply the EGNN model to semiconducting polymers consisting of hundreds of atoms, which are difficult to calculate by rigorous correlated methods such as CCSD(T). Semiconducting polymers are organic macromolecules with small energy gap and high electrical conductivity compared to insulating polymers. Because of these electronic features, semiconducting polymers attract broad research interests both for the fundamental understanding of quantum chemistry [49] and for applications in semiconductor industry [52]. The essential electronic properties of semiconducting polymers originates from the $\pi$-bonds with delocalized molecular orbitals. As the delocalized molecular orbitals extend through the whole macroscopic molecule (Fig. 4a), the polymers’ electronic properties also involve long-range correlation, making it challenging for ML methods. Therefore, it is important to examine whether the EGNN model can capture semiconducting polymers’ electronic properties involving delocalized molecular orbitals.
@@
Three kinds of semiconducting polymers, trans-polyacetylene (t-PA), cyclic polyacetylene (c-PA), and polyphenylene (PPP), are studied using our EGNN model. The model correctly captures the delocalized $\pi$-bond feature of frontier orbitals (highest occupied molecular orbital (HOMO) and lowest unoccupied molecular orbital (LUMO)), as shown in Fig. 4a. In the t-PA, the HOMO and LUMO are on the carbon-carbon double and single bonds, respectively, consistent with the predictions from the renowned Su–Schrieffer–Heeger (SSH) model [49]. In the c-PA, the left/right mirror symmetry of the molecule is correctly reflected in the predicted frontier orbitals, which exhibits odd parity upon mirror reflection. Both the HOMO and LUMO have two anti-phase boundaries where the wavefunctions shift from double bonds to single bonds. The same symmetry is also reflected in the PPP with odd-parity HOMO and even-parity LUMO. It should be emphasized that the molecular orbitals are not included as fitting targets in our model training. However, as the model learned relevant properties such as electric moments, atomic charge, and bond order, it gives at least qualitatively correct predictions for frontier orbitals in these semiconducting polymers.
@@
Various important electronic properties of semiconducting polymers depend on the chain length, including the energy gap $E_{\rm g}$ and polarizability $\alpha$. We calculate such chain-length dependence (up to more than 400 atoms) using our EGNN method, as shown in Fig. 4b,c. One can see that $E_{\rm g}$ is larger for oligomers with shorter chain length due to the quantum confinement effect and converges to a smaller value for long polymer. The converged energy gap for long t-PA and PPP polymers calculated by our EGNN are in reasonable agreement with the experimental values [49, 48] (relative errors within 10%), which are shown as squares in Fig. 4b. The longitudinal static electric polarizability $\alpha_{xx}$ (per monomer) is positively related to the polymer chain length. This is because in longer chains, more delocalized electron distributions can have larger displacements under external electric field. The predicted $\alpha_{xx}$ for t-PA oligomers and PPP polymers are in perfect agreement with previous correlated calculations using the high-accuracy MP2 method [50, 51], as shown in Fig. 4c. The chain length-dependent $E_{\rm g}$ and $\alpha$ of cyclic PA, to the best of our knowledge, have not been reported. We provide their values as a prediction to be examined by future work.
@@ III Conclusion and Outlook
In this work, we developed a multi-task learning method for predicting molecular electronic structure with coupled-cluster-level accuracy. In our method, an EGNN is trained on a series of molecular properties in order to capture their shared underlying representation, the effective electronic Hamiltonian. Our EGNN model shows significantly lower computational costs and higher prediction accuracy on various molecular properties than the widely-used B3LYP functional.
Compared to other works that take the electronic structure as a direct fitting target [19, 27, 25, 26], our method takes the electronic structure as a shared representation to predict various molecular properties. This approach, on the one hand, enables training on data beyond the DFT accuracy; on the other hand, relieve the burden of fitting all matrix elements of the electronic Hamiltonian, which are not direct quantum observable. The later enables us to go beyond the minimal basis used in previous works by a relatively small NN with only 511,589 parameters. With the electronic structure as a physics-informed representation, our method exhibits generalization capability for challenging systems with delocalized molecular orbitals. In such systems, atomic structure at one position has long-range influence to electronic features (such as electron density) at another distant position. Although the direct output of our EGNN, the matrix elements of $\bf{V}_{\theta}$, only depends on local atomic environments, the long-range influence is captured through the output-layer calculations based on the rules of quantum mechanics. In comparison, such long-range influence is unlikely to be captured by directly using graph neural network or local atomic descriptors to predict molecular properties without a physics-informed representation.
Although the current EGNN model is limited to hydrocarbon molecules, our method can be readily applied to systems with different elements. Applying the method to datasets with more elements can produce a general-purpose electronic structure predictor, which is left to future work.
@@ IV.1 Graph encoding of atomic configuration
As the hydrocarbon molecules we study are all close-shell molecules, we use spin-restricted DFT calculations to obtain ${\bf F}$ and we assume ${\bf V}_{\theta}$ is spin-independent throughout this paper. Namely, the spin-up and spin-down molecular orbitals and energy levels are the same, and all molecular orbitals are either doubly occupied or vacant. The electronic energy of the BP86 DFT calculation $E_{\rm BP86}$ equals the band structure energy $2\sum_{i=1}^{n_{e}/2}\epsilon_{i}$ (where $\epsilon_{i}$ is the $i$-th molecular orbital energy level and $n_{e}$ is the number of electrons) plus a many-body energy $E_{\rm MB}$:
@@
where ${\rm eig}_{i}$ is a function that returns the $i$-th lowest eigenvalue of a matrix, and we use the fact that the energy level $\epsilon_{i}$ is the eigenvalue of the Lowdin-symmetrized Hamiltonian ${\rm eig}_{i}(S^{-1/2}FS^{-1/2})$ [35]. After this transformation, the KS effective Hamiltonian (both before and after the $\bf{V}_{\theta}$ correction) already includes the many-body energy, and the total electronic energy is just the summation of band structure energies. Adding the $E_{\rm MB}$ term does not change the eigenfunction and relative energy levels, so that all other properties are not changed by adding the $E_{\rm MB}$ term.
@@ IV.2 Architecture of the convolutional layer
In the following technical description, we use terminologies defined in the e3nn documentation. One can refer to Ref. [30] for further information. In the convolutional layer, the input feature first goes through a $N_{\rm species}\times N_{\rm species}$ linear transformation (the first Self-Interaction block, $N_{\rm species}$ is the number of different elements in the system) and an activation layer (the first non-Linearity block). All activation layers in our EGNN are realized by tanh function acting on scalar features.
@@
Finally, the output features are used to construct the correction matrix $\bf{V}_{\theta}$ through Eq. (2). $V_{\rm node}(\mathbf{x}_{I,\rm out})$ first apply a linear layer from input dimension of Irreps2 to output dimension $\text{Irreps}(I)^{\otimes 2}$, where:
@@
The output dimension corresponds to the irreducible representation of the block diagonal terms of the Hamiltonian (as cc-pVDZ basis of hydrogen includes two s orbitals (0e) and one group of p orbitals (1o); while that of carbon includes three s orbitals (0e), two groups of p orbitals (1o), and one group of $d$ orbitals (2e)). The output is then arranged into the $N_{I}\times N_{I}$ matrix form, $V_{I,\rm out}$, according to the Wigner-Eckart theorem [26], and symmetrized to obtain $V_{\rm node}(\mathbf{x}_{I,\rm out})=\frac{\lambda_{V}}{2}(V_{I,\rm out}+V_{I,\rm out}^{T})$. $\lambda$ is a constant hyperparameter set as 0.2 for our model. Similarly, the off-diagonal term $V_{\rm edge}(\mathbf{f}_{IJ,\rm out})$ in Eq. (2) applies a linear layer from input dimension of Irreps2 to output dimension Irreps($I,J$) that equals: